i am 63 yrs old,many years out of school,can you give me a basic formula for calculating percentages,i have a test on wednesday,remember i am old,have not kept up with math
thanks
example 23% of 513

Answer by Alan3354(69443)

example 23% of 513
-------------------
% means part of 100
23% = 23/100 = 0.23
0.23*513 = 117.99